OpenStudy (anonymous):

Solve: (x - 8)(x + 11) = 0. Answer {8, -11} {-8, 11} {0, 11/8} no solution

OpenStudy (anonymous):

(x - 8)(x + 11) = 0. i..e (x - 8)=0 or (x + 11) = 0. i..e x = 8 or(x =- 11 therefore solutin set is x={8, -11}

OpenStudy (anonymous):

don't say! u already knew the answer again @charlieberzak

OpenStudy (anonymous):

idk you at charlie but ur pretty cute do u have twitter

OpenStudy (anonymous):

x=8 and -11

OpenStudy (anonymous):

I wanted to be sure @kritima and thank you @blahkatherine ^^ no i don't have twitter D: